What is the total number of molecules in 13.41L of O2?
1
Expert's answer
2017-11-29T12:58:07-0500
n(0₂)=13.41L/22.4L/Mol=0.6Mol
N(O₂)=0.6*6.02*10²³=3.612*10²³